OpenStudy (decarr432):

How do I find the slope and y intercept What is the slope and y-intercept of the equation 3(y − 2) + 6(x + 1) − 2 = 0

OpenStudy (decarr432):

@pooja195 you know how to solve?

pooja195 (pooja195):

We need to solve for y first... Any ideas on where to start?

OpenStudy (decarr432):

get y alone?

pooja195 (pooja195):

Yes... 3(y − 2) + 6(x + 1) − 2 = 0 Start by distributing \[\huge~\rm~\bf~\color{green}{3}\color{red}{(y − 2)}+\color{green}{ 6}\color{red}{(x + 1)}− 2 = 0 \]

OpenStudy (decarr432):

(3y-6)+(6x+6)-2=0? like so?

pooja195 (pooja195):

\[\huge~\rm~\bf~3y-6+6x+6-2=0\] combine like terms

OpenStudy (decarr432):

okay so 3y+6x-2

pooja195 (pooja195):

ok good add 2 to both sides \[\huge~\rm~\bf~ 3y+6x-2\color{red}{+2} =0 \color{red}{+2} \]

pooja195 (pooja195):

what equation do we have now?

OpenStudy (decarr432):

3y+6x=2?

pooja195 (pooja195):

Good 3y+6x=2 To get y alone what would you do next?

OpenStudy (decarr432):

divide by 3?

pooja195 (pooja195):

mmm not yet...try again

OpenStudy (decarr432):

Subtract 6x?

pooja195 (pooja195):

Perfect what equation do we get once we do that?

OpenStudy (decarr432):

3y=2-6x

pooja195 (pooja195):

Good let's organize that in slope intercept form though y=mx+b \[\huge~\rm~ \bf~3y=-6x+2\] What's next?

OpenStudy (decarr432):

divide by 3

pooja195 (pooja195):

good what equation do we get once we do that?

OpenStudy (decarr432):

y=-6x+2/3?

pooja195 (pooja195):

hmm let's try that again, Good let's organize that in slope intercept form though \[\huge~\rm~\bf~ \frac{ 3y=-6x+2 }{ \color{purple}{3~~~~~~~~~~3~~~~~3} }\]

pooja195 (pooja195):

3/3=?

OpenStudy (decarr432):

1

pooja195 (pooja195):

good so we have y=.... -6/3=?

OpenStudy (decarr432):

Sorry I just got really confused

pooja195 (pooja195):

It's alright we're here to learn! :) What are you confused on?

OpenStudy (decarr432):

good so we have y=.... -6/3=?

OpenStudy (decarr432):

That

pooja195 (pooja195):

\(\color{#0cbb34}{\text{Originally Posted by}}\) @pooja195 hmm let's try that again, Good let's organize that in slope intercept form though \[\huge~\rm~\bf~ \frac{ 3y=-6x+2 }{ \color{purple}{3~~~~~~~~~~3~~~~~3} }\] \(\color{#0cbb34}{\text{End of Quote}}\) Remeber that we are simply dividing everything by 3 to isolate the y

OpenStudy (decarr432):

Yeah but how do we divide 2 by 3??

pooja195 (pooja195):

You don't it remains the same :-)

pooja195 (pooja195):

-6/3=?

OpenStudy (decarr432):

Ah okay

pooja195 (pooja195):

You'll see at the end, what i mean.

OpenStudy (decarr432):

Thanks for the help btw

pooja195 (pooja195):

yw ^.^ Let's finish up the problem!

OpenStudy (decarr432):

Okay

pooja195 (pooja195):

Can you tell me what equation we would get when we divide all the numbers by 3?

OpenStudy (decarr432):

y=(-6x+2)/3

pooja195 (pooja195):

almost.... You need to divide EACH number by 3 -6/3=?

OpenStudy (decarr432):

okay but what about 2?

pooja195 (pooja195):

worry about that later answer what im asking.

OpenStudy (decarr432):

-2

pooja195 (pooja195):

Perfect so y=-2x+___ 2/3=2/3 it's already in its simplest form therefore it remains the same y=-2x+2/3 y=mx+b m=slope b=y intercept m=? b=?

OpenStudy (decarr432):

m=2 b=2/3 right?

pooja195 (pooja195):

*m=-2 <--don't forget the - sign b=2/3 <---correct

OpenStudy (decarr432):

Oh okay so my slope is -2 and y intercept is positive 2/3

pooja195 (pooja195):

yes ^_^

OpenStudy (decarr432):

Okay thank you so much

pooja195 (pooja195):

yw ^_^